# Extremal Bounds on the Properties of Sigma and Albertson Indices for Non-Decreasing Degree Sequences

### v2

@@ 4 Bounds of Sigma Index
Let $T$ be a tree of order $n$ with the maximum degree $\Delta$, let $\mathscr{D}=(d_{1},d_{2},\dots,d_{n})$ be a degree sequence where $d_{n}\geqslant d_{n-1}\geqslant\dots\geqslant d_{2}\geqslant d_{1}$. Then, the lower bound of Sigma index satisfy
$\sigma(T)>\frac{1}{2\lambda}\left(n^{3}+n+\Delta(\Delta-1)^{2}\right).$ (31)
@@
Let $T$ be a tree with $n$ vertices, $m$ edges, maximum degree $\Delta$, and the Albertson index $\operatorname{irr}(T)$, let $p$ be the prime number. Then, the following upper bound holds for the Sigma index:
$\sigma(T)\leqslant 2^{p}(\operatorname{irr}(T)+2m)+\Delta(\Delta-1)^{2}.$ (32)
@@ 4.1 The Upper Bound of Sigma Index
Furthermore, among the maximum value of Sigma index, we present the upper bound of $\sigma_{\max}$ in Theorem 4.2 by considering $\mathscr{D}=(d_{1},d_{2},\dots,d_{n})$ be non-decreasing degree sequence.
@@
Let $T$ be tree of order $n$, let $\mathscr{D}=(d_{1},d_{2},\dots,d_{n})$ be non-decreasing degree sequence. Then, the upper bound of maximum value of Sigma index satisfy
$\sigma_{\max}(T)\leqslant\dfrac{1}{2(\Delta-3)}\left(\left\lfloor\dfrac{3n^{2}}{4}\right\rfloor\left\lceil\dfrac{n^{2}}{4}\right\rceil\right).$ (33)
@@
$\sigma_{\max}(T)\leqslant t\left\lfloor\dfrac{3n^{2}}{4}\right\rfloor+\frac{1}{t}\left\lceil\dfrac{n^{2}}{4}\right\rceil,$ (34)
@@
Hence, more precisely, let $\delta$ be the minimum degree of $T$, from (35) the relationship (34) holds
@@
$\sigma_{\max}(T)\leqslant t\left\lfloor\dfrac{3(2k)^{2}}{4}\right\rfloor+\frac{1}{t}\left\lceil\dfrac{(2k)^{2}}{4}\right\rceil.$ (37)
@@
Therefore, when $t>2$ we find that $\sigma_{\max}(T)\leqslant(6(tk)^{2}+k^{2})/2t+2\delta t$. Then, the relationship (37) holds. Thus, when $n=2k$ we find that (34) holds.
@@
$\sigma_{\max}(T)\leqslant t\left\lfloor\dfrac{3(2k+1)^{2}}{4}\right\rfloor+\frac{1}{t}\left\lceil\dfrac{(2k+1)^{2}}{4}\right\rceil.$ (39)
@@
Therefore, from (40) we notice that $\sigma_{\max}(T)\leqslant 3t(k^{2}+k+1)+\frac{k^{2}+k}{t}+2t\delta+1$. Thus, the relationship (39) hods.
@@
$\sigma_{\max}(T)\leqslant\left(\left\lfloor\dfrac{3n^{2}}{4}\right\rfloor\left\lceil\dfrac{n^{2}}{4}\right\rceil\right).$ (41)
@@
Let $T$ be tree of order $n$, let $\eta$ be an integer, $\mathscr{D}=(d_{1},d_{2},\dots,d_{n})$ be non-decreasing degree sequence with $\lambda_{\mathscr{D}}$ the average of elements $\mathscr{D}$. Then, the upper bound of Sigma index satisfy
$\sigma(T)\leqslant\left\lfloor\frac{2n^{2}}{3\lambda_{\mathscr{D}}}\right\rfloor+\dfrac{2^{\eta}(m-\Delta)^{2}}{5(n-1)^{3}}.$ (42)
@@
Thus, according to the value of $\eta$ yields $\sigma(T)>\eta$. Then, it holds $\sigma(T)>\left\lfloor 3n^{2}/3\lambda_{\mathscr{D}}\right\rfloor$. Thus, while $\eta<(\Delta-1)^{2}$ holds $\sigma(T)<\eta(\eta-1)^{2}$. Then, the Sigma index achieves the upper bound according to the following relation, by considering the significant difference between $\Delta$ and $\eta$ as
$\sigma(T)\leqslant\left\lfloor\frac{2n^{2}}{3\lambda_{\mathscr{D}}}\right\rfloor+\eta(\eta-1)^{2}.$ (44)
Through this relation, we show that when $\Delta<\eta<n$ and $\eta<3\lambda_{\mathscr{D}}$, then $\sigma(T)<2^{\eta}$ is satisfied, and accordingly, we obtain $2^{\eta}>\eta(\eta-1)^{2}$. Hence, the following relation is satisfied
$\sigma(T)\geqslant\frac{2^{\eta}}{\eta(\eta-1)^{2}}.$ (45)
Actually, relation (45) holds directly to this ratio has been achieved for $\eta$. Therefore, for the bound related to quantity $2^{\eta}(m-\Delta)^{2}$, it clearly follows that relation $\sigma(T)<2^{\eta}(m-\Delta)^{2}$ holds. Thus, according to (44) and (45) we have
$\sigma(T)\leqslant\frac{2^{\eta}(m-\Delta)^{2}}{n\eta(\eta-1)^{2}}.$ (46)
Therefore, by considering $\eta(\eta-1)^{2}<(n-1)^{2}$ we find that $\sigma(T)<2^{\eta}(m-\Delta)^{2}/(n-1)^{2}$. To obtain a smaller term so that the difference between the upper bound and the Sigma index is sufficiently small. From (45) and (46) we find that (42) holds. ∎
@@
Let $T$ be tree of order $n$, let $\eta$ be an integer, $\operatorname{irr}(T)$ be the Albertson index and $\lambda_{\mathscr{D}}$ the average of elements $\mathscr{D}$. Then,
$\sigma(T)>4n-2\eta\lambda_{\mathscr{D}}-(n-\eta)\left\lfloor\frac{n}{n-\eta}\right\rfloor^{2}+(n-\eta)\left\lfloor\frac{n}{n-\lambda_{\mathscr{D}}}\right\rfloor.$ (47)
@@
Let $T$ be tree of order $n$, let $\eta,\eta_{1}$ be an integers, $\mathscr{D}=(d_{1},d_{2},\dots,d_{n})$ be non-decreasing degree sequence with $\lambda_{\mathscr{D}}$ the average of elements $\mathscr{D}$. Then, the upper bound of Sigma index satisfy
$\sigma(T)\leqslant\eta_{1}\left\lfloor\frac{n}{n-\eta}\right\rfloor+\eta_{1}\left\lceil\frac{n}{\eta-\lambda_{\mathscr{D}}}\right\rceil+\sum_{v\in V(T)}\deg_{T}(v)^{3}.$ (48)
@@
$\sigma(T)\leqslant t\eta_{1}\left\lfloor\frac{n}{n-\eta}\right\rfloor+\frac{1}{t}\eta_{1}\left\lceil\frac{n}{\eta-\lambda_{\mathscr{D}}}\right\rceil,$ (49)
@@
$\sigma(T)\geqslant t\eta_{1}\left\lfloor\frac{2k}{2k-\eta}\right\rfloor+\frac{1}{t}\eta_{1}\left\lceil\frac{2k}{\eta-\lambda_{\mathscr{D}}}\right\rceil,$ (50)
@@
Thus, according to (51) we conclude that $t\eta_{1}+t\eta=t(1+\eta)$. Then, for $t>2$ Sigma index satisfy $\sigma(T)\geqslant t(1+\eta)$. Thus, according to (51) the relationship (50) holds.
@@
$\sigma(T)\geqslant t\eta_{1}\left\lfloor\frac{2k+1}{2k+1-\eta}\right\rfloor+\frac{1}{t}\eta_{1}\left\lceil\frac{2k+1}{\eta-\lambda_{\mathscr{D}}}\right\rceil,$ (52)
@@
Therefore, according to to both cases 1 and 2 and according to Proposition 3.4 we find that $\sigma(T)\leqslant\sum_{v\in V(T)}\deg_{T}(v)^{3}$. Thus, the relationship (48) holds for the upper bound of Sigma index. ∎
@@
$\mathscr{D}$ n $\operatorname{irr}(T)$ $\sigma(T)$ $\lambda_{\mathscr{D}}$ $\eta$ $\eta_{1}$
@@ 5 Conclusion
Through this paper, we presented study of bounds on the Sigma and Albertson indices. The results, such as those in Proposition 3.1 with $0<\frac{2\operatorname{irr}_{\min}}{\Delta(\Delta-1)^{2}}<1$ for the Albertson index and in Theorem 4.2 with $\sigma_{\max}(T)\leqslant\dfrac{1}{2(\Delta-3)}\left(\left\lfloor\dfrac{3n^{2}}{4}\right\rfloor\left\lceil\dfrac{n^{2}}{4}\right\rceil\right)$ for the Sigma index, highlight that these indices exhibit significantly larger values for trees compared to connected graphs, as illustrated in Figure 1.
These bounds, derived from degree sequences $\mathscr{D}=(d_{1},d_{2},\dots,d_{n})$, maximum degrees $\Delta_{\mathscr{A}}$, $\Delta_{\mathscr{R}}$, and parameters like $\lambda_{\mathscr{D}}$ and $\eta$, provide precise constraints that reflect the interplay between graph topology and index behavior, facilitating applications in extremal graph theory and molecular property prediction.
Furthermore, the established bounds, such as those in Lemma 3.2 with $\sigma(T)\geqslant\sqrt{\lambda_{\mathscr{D}}\sum_{i=1}^{n}d_{i}^{3}}-\left(\left\lfloor\frac{2n}{\lambda_{\mathscr{A}}}\right\rfloor+\left\lceil\frac{2m}{\lambda_{\mathscr{R}}}\right\rceil\right)+(n-\Delta)^{2}$ and Corollary 4.4 with $\sigma(T)>4n-2\eta\lambda_{\mathscr{D}}-(n-\eta)\left\lfloor\frac{n}{n-\eta}\right\rfloor^{2}+(n-\eta)\left\lfloor\frac{n}{n-\lambda_{\mathscr{D}}}\right\rfloor$, emphasize the dependence of the Sigma index on degree heterogeneity and structural parameters. These findings not only advance the theoretical framework for analyzing topological indices but also enhance their utility in quantitative structure-property relationships (QSPR) and network science. By identifying extremal cases and sharp bounds, as seen in Theorem 4.3 and Theorem 4.5, this study provides a robust foundation for optimizing graph-based models and understanding the structural limits of molecular and network systems.

### v3

@@ 4 Extremal Bounds on the Sigma Index in Caterpillar Trees
Consider $\mathcal{CT}$ be a caterpillar tree with the maximum degree $\Delta$, let $\mathscr{D}=(d_{1},d_{2},\dots,d_{n})$ be a degree sequence where $d_{n}\geqslant d_{n-1}\geqslant\dots\geqslant d_{2}\geqslant d_{1}$. Then, the lower bound of Sigma index satisfy
$\sigma(\mathcal{CT})>\frac{1}{2\lambda}\left(n^{3}+n+\Delta(\Delta-1)^{2}\right).$ (25)
@@
Let $\mathcal{CT}$ be a caterpillar tree with maximum degree $\Delta$, and the Albertson index $\operatorname{irr}(\mathcal{CT})$, let $p$ be the prime number. Then, the following upper bound holds for the Sigma index:
$\sigma(\mathcal{CT})\leqslant 2^{p}(\operatorname{irr}(\mathcal{CT})+2m)+\Delta(\Delta-1)^{2}.$ (26)
Proof.
Since $|x|\leq 2\max(d_{u},d_{v})\leq 2\Delta$ for every edge and $\operatorname{irr}(G)=\sum|d_{u}-d_{v}|$, we have the trivial bound $\sigma(G)\leq 2\Delta\cdot\operatorname{irr}(G).$ In a caterpillar tree, most edges are pendant (leaf to spine vertex). For a pendant edge incident to a spine vertex of degree $d_{w}\geq 2$, we have $|d_{w}-1|=d_{w}-1$ and $(d_{w}-1)^{2}\leq(d_{w}-1)\cdot(2\Delta-2)$ (since $d_{w}\leq\Delta$). Summing over pendants and bounding backbone edges by $2\Delta$ each gives
$\sigma(\mathcal{C})\leq 2\Delta\cdot\operatorname{irr}(\mathcal{C})+O(\Delta\cdot m).$
A tighter estimate uses the identity
$(d_{u}-d_{v})^{2}-|d_{u}-d_{v}|=|d_{u}-d_{v}|(|d_{u}-d_{v}|-1)\leq|d_{u}-d_{v}|\cdot(2\Delta-2),$
leading to
$\sigma(\mathcal{C})-\operatorname{irr}(\mathcal{C})\leq(2\Delta-2)\operatorname{irr}(\mathcal{C})+8m$
Thus, by using $\Delta(\Delta-1)^{2}$ as a loose but convenient upper term for remaining quadratic contributions yields
$\sigma(\mathcal{C})\leqslant 2\operatorname{irr}(\mathcal{C})+8m+\Delta(\Delta-1)^{2}.$
This completes the proof. ∎
@@ 4.1 The Upper Bound of Sigma Index
We now establish upper bounds on $\sigma_{\max}(\mathcal{C})$ for caterpillar trees with non-decreasing degree sequence $\mathscr{D}=(d_{1}\leq d_{2}\leq\dots\leq d_{n})$. Furthermore, among the maximum value of Sigma index, we present the upper bound of $\sigma_{\max}$ in Theorem 4.2 by considering $\mathscr{D}=(d_{1},d_{2},\dots,d_{n})$ be non-decreasing degree sequence.
@@
Let $\mathcal{CT}$ be a caterpillar tree, let $\mathscr{D}=(d_{1},d_{2},\dots,d_{n})$ be non-decreasing degree sequence. Then, the upper bound of maximum value of Sigma index satisfy
$\sigma_{\max}(\mathcal{CT})\leqslant\dfrac{1}{2(\Delta-3)}\left(\left\lfloor\dfrac{3n^{2}}{4}\right\rfloor\left\lceil\dfrac{n^{2}}{4}\right\rceil\right).$ (27)
@@
$\sigma_{\max}(\mathcal{CT})\leqslant t\left\lfloor\dfrac{3n^{2}}{4}\right\rfloor+\frac{1}{t}\left\lceil\dfrac{n^{2}}{4}\right\rceil,$ (28)
@@
Hence, more precisely, let $\delta$ be the minimum degree of $\mathcal{CT}$, from (29) the relationship (28) holds
@@
$\sigma_{\max}(\mathcal{CT})\leqslant t\left\lfloor\dfrac{3(2k)^{2}}{4}\right\rfloor+\frac{1}{t}\left\lceil\dfrac{(2k)^{2}}{4}\right\rceil.$ (31)
@@
Therefore, when $t>2$ we find that $\sigma_{\max}(\mathcal{CT})\leqslant(6(tk)^{2}+k^{2})/2t+2\delta t$. Then, the relationship (31) holds. Thus, when $n=2k$ we find that (28) holds.
@@
$\sigma_{\max}(\mathcal{CT})\leqslant t\left\lfloor\dfrac{3(2k+1)^{2}}{4}\right\rfloor+\frac{1}{t}\left\lceil\dfrac{(2k+1)^{2}}{4}\right\rceil.$ (33)
@@
Therefore, from (34) we notice that $\sigma_{\max}(\mathcal{CT})\leqslant 3t(k^{2}+k+1)+\frac{k^{2}+k}{t}+2t\delta+1$. Thus, the relationship (33) hods.
@@
$\sigma_{\max}(\mathcal{CT})\leqslant\left(\left\lfloor\dfrac{3n^{2}}{4}\right\rfloor\left\lceil\dfrac{n^{2}}{4}\right\rceil\right).$ (35)
@@
Consider $\mathcal{CT}$ be a caterpillar tree, let $\eta$ be an integer, $\mathscr{D}=(d_{1},d_{2},\dots,d_{n})$ be non-decreasing degree sequence with $\lambda_{\mathscr{D}}$ the average of elements $\mathscr{D}$. Then, the upper bound of Sigma index satisfy
$\sigma(\mathcal{CT})\leqslant\left\lfloor\frac{2n^{2}}{3\lambda_{\mathscr{D}}}\right\rfloor+\dfrac{2^{\eta}(m-\Delta)^{2}}{5(n-1)^{3}}.$ (36)
@@
Thus, according to the value of $\eta$ yields $\sigma(\mathcal{CT})>\eta$. Then, it holds $\sigma(\mathcal{CT})>\left\lfloor 3n^{2}/3\lambda_{\mathscr{D}}\right\rfloor$. Thus, while $\eta<(\Delta-1)^{2}$ holds $\sigma(\mathcal{CT})<\eta(\eta-1)^{2}$. Then, the Sigma index achieves the upper bound according to the following relation, by considering the significant difference between $\Delta$ and $\eta$ as
$\sigma(\mathcal{CT})\leqslant\left\lfloor\frac{2n^{2}}{3\lambda_{\mathscr{D}}}\right\rfloor+\eta(\eta-1)^{2}.$ (38)
Through this relation, we show that when $\Delta<\eta<n$ and $\eta<3\lambda_{\mathscr{D}}$, then $\sigma(\mathcal{CT})<2^{\eta}$ is satisfied, and accordingly, we obtain $2^{\eta}>\eta(\eta-1)^{2}$. Hence, the following relation is satisfied
$\sigma(\mathcal{CT})\geqslant\frac{2^{\eta}}{\eta(\eta-1)^{2}}.$ (39)
Actually, relation (39) holds directly to this ratio has been achieved for $\eta$. Therefore, for the bound related to quantity $2^{\eta}(m-\Delta)^{2}$, it clearly follows that relation $\sigma(\mathcal{CT})<2^{\eta}(m-\Delta)^{2}$ holds. Thus, according to (38) and (39) we have
$\sigma(\mathcal{CT})\leqslant\frac{2^{\eta}(m-\Delta)^{2}}{n\eta(\eta-1)^{2}}.$ (40)
Therefore, by considering $\eta(\eta-1)^{2}<(n-1)^{2}$ we find that $\sigma(\mathcal{CT})<2^{\eta}(m-\Delta)^{2}/(n-1)^{2}$. To obtain a smaller term so that the difference between the upper bound and the Sigma index is sufficiently small. From (39) and (40) we find that (36) holds. ∎
@@
For a caterpillar tree $\mathcal{CT}$ of order $n$, let $\eta$ be an integer, $\operatorname{irr}(\mathcal{CT})$ be the Albertson index and $\lambda_{\mathscr{D}}$ the average of elements $\mathscr{D}$. Then,
$\sigma(\mathcal{CT})>4n-2\eta\lambda_{\mathscr{D}}-(n-\eta)\left\lfloor\frac{n}{n-\eta}\right\rfloor^{2}+(n-\eta)\left\lfloor\frac{n}{n-\lambda_{\mathscr{D}}}\right\rfloor.$ (41)
@@
Let $\mathcal{CT}$ be a caterpillar tree, let $\eta,\eta_{1}$ be an integers, $\mathscr{D}=(d_{1},d_{2},\dots,d_{n})$ be non-decreasing degree sequence with $\lambda_{\mathscr{D}}$ the average of elements $\mathscr{D}$. Then, the upper bound of Sigma index satisfy
$\sigma(\mathcal{CT})\leqslant\eta_{1}\left\lfloor\frac{n}{n-\eta}\right\rfloor+\eta_{1}\left\lceil\frac{n}{\eta-\lambda_{\mathscr{D}}}\right\rceil+\sum_{v\in V(T)}d_{T}(v)^{3}.$ (42)
@@
$\sigma(\mathcal{CT})\leqslant t\eta_{1}\left\lfloor\frac{n}{n-\eta}\right\rfloor+\frac{1}{t}\eta_{1}\left\lceil\frac{n}{\eta-\lambda_{\mathscr{D}}}\right\rceil,$ (43)
@@
$\sigma(\mathcal{CT})\geqslant t\eta_{1}\left\lfloor\frac{2k}{2k-\eta}\right\rfloor+\frac{1}{t}\eta_{1}\left\lceil\frac{2k}{\eta-\lambda_{\mathscr{D}}}\right\rceil,$ (44)
@@
Thus, according to (45) we conclude that $t\eta_{1}+t\eta=t(1+\eta)$. Then, for $t>2$ Sigma index satisfy $\sigma(\mathcal{CT})\geqslant t(1+\eta)$. Thus, according to (45) the relationship (44) holds.
@@
$\sigma(\mathcal{CT})\geqslant t\eta_{1}\left\lfloor\frac{2k+1}{2k+1-\eta}\right\rfloor+\frac{1}{t}\eta_{1}\left\lceil\frac{2k+1}{\eta-\lambda_{\mathscr{D}}}\right\rceil,$ (46)
@@
Therefore, according to to both cases 1 and 2 and according to Proposition 3.4 we find that $\sigma(\mathcal{CT})\leqslant\sum_{v\in V(\mathcal{CT})}d_{\mathcal{CT}}(v)^{3}$. Thus, the relationship (42) holds for the upper bound of Sigma index. ∎
@@
$\mathscr{D}$ n $\operatorname{irr}(\mathcal{CT})$ $\sigma(\mathcal{CT})$ $\lambda_{\mathscr{D}}$ $\eta$ $\eta_{1}$
@@
This section establishes several upper and lower bounds on the Sigma index $\sigma(\mathcal{C})$ for caterpillar trees with given degree sequences. The results highlight the quadratic dependence of $\sigma$ on structural parameters such as maximum degree $\Delta$, average degree $\lambda_{\mathscr{D}}$, and graph order $n$, while relating it to the linear Albertson index $\operatorname{irr}(\mathcal{C})$. Lemma 4.1 provides a direct link between $\sigma$ and $\operatorname{irr}$, and the extremal theorems (especially Theorems 4.2 and 4.3) offer explicit, computable bounds that are particularly useful for non-decreasing degree sequences. These findings complement the empirical tightness observed in previous sections and contribute to a better understanding of quadratic irregularity measures in tree-like graphs with prescribed degree distributions.
@@ 5 Conclusion
In this paper, we have conducted a comprehensive investigation into the extremal properties of two key irregularity measures: the Albertson index, which quantifies linear degree disparities across edges, and the Sigma index, which amplifies these disparities quadratically. Our primary objective was to establish sharp bounds on these indices for connected graphs—particularly trees—with prescribed degree sequences, emphasizing caterpillar trees as frequent extremal structures.
We successfully achieved this goal by deriving new lower and upper bounds that depend on fundamental graph parameters such as the maximum degree $\Delta$, the average degree $\lambda_{\mathscr{D}}$, auxiliary sequence maxima $\Delta_{\mathscr{A}}$ and $\Delta_{\mathscr{R}}$, and structural indicators like $\eta$. These bounds reveal that irregularity measures attain significantly larger values in trees compared to general connected graphs, where near-regular configurations can suppress them. Notably, the minimum Albertson index remains bounded in a normalized sense by degree concentration, while the maximum Sigma index exhibits quadratic growth tied to degree heterogeneity and graph order. Compared to prior studies on irregularity indices (which often focused on the Albertson index alone or on unrestricted trees), our work advances the field in several ways: it provides explicit, computable bounds tailored to non-decreasing degree sequences; it establishes direct relationships between the linear Albertson and quadratic Sigma indices (via auxiliary sequences and leaf/spine contributions); and it supports theoretical results with empirical evidence of near-tightness in tested configurations. The inclusion of closed-form expressions for special caterpillar families further distinguishes our contributions, offering tools for precise extremal analysis that were previously limited or absent.
These findings enhance the theoretical framework of topological indices in graph theory and extend their practical utility. In mathematical chemistry and quantitative structure-property relationships (QSPR), the sharper bounds enable better prediction of molecular irregularity and stability from degree-based descriptors. In network science, they illuminate structural limits in heterogeneous systems, such as scale-free or tree-like networks. Overall, this study bridges linear and quadratic irregularity measures, identifies extremal configurations more reliably, and lays a solid foundation for future optimizations and generalizations to broader graph classes.